# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v2

@@ 2.2 A closer look at the internals of the ME engine
In the production version of the CUDACPP plugin before any kernel splitting changes (hereafter “ihel0”), v1.00.02 in MG5aMC v3.6.3, the ME computational engine on GPUs is essentially a single and very large monolithic GPU kernel, named sigmaKin. This performs all of the relevant operations that are needed to compute the output ME from the input momenta for each event (one notable exception being the calculation of other couplings that depend on $\alpha_{s}$, which is delegated to a separate kernel). Very schematically, this is represented in the pseudo-code in Table 1. In particular, sigmaKin performs the following operations for each event: it keeps a running sum of the ME over helicities; it loops on all possible helicities of the external particles (or, more precisely, only on a pre-determined set of “good” helicities whose contribution is non-zero); it adds to the running sum of the ME the contribution from each helicity, by calling a device function calculate_wavefunctions which internally involves for each QCD color flow the calculation of particle and propagator wavefunctions and that of dual amplitudes, followed by the sum over all color flows using a quadratic form based on a color matrix; after the end of the helicity loop, sigmaKin randomly draws a helicity and a color for the generated event based on the ME contributions from individual helicities and colors. The calculation of helicity amplitudes for Feynman diagrams, in particular, is described in detail in Ref. bib:vchep2021 (see Fig. 1 therein); complementary details are also given in Ref. bib:zw2025 . As shown schematically in Table 1, this involves three types of elements to compute: the wavefunctions of external (initial/final state) particles, via functions like VXXXXX; the wavefunctions of internal propagators, e.g. via VVV1P0_1; the dual amplitudes for a given colour flow (“jamps”), e.g. via FFV1_0. The names and roles of these functions in CUDACPP are the same as in the original HELAS bib:helas and ALOHA bib:aloha implementations.
@@
    Memory layout and data access. The specific layout chosen for storing data inside the raw memory buffers allocated in the previous step is encapsulated in a separate set of memory access classes. All these classes have methods named kernelAccess, or some variation of this. On the GPU, these methods take as input the fptype* pointer associated to a raw memory buffer for nevt events, and they return as output a data item for the single event ievt processed by the GPU thread where the code is executed, typically indexed by the identifier of the GPU thread, blockDim.x*blockIdx.x+threadIdx.x. In some cases, these methods have additional parameters: method kernelAccessIp4IparConst in the DeviceAccessMomenta class, for instance, takes two additional parameters ip4 and ipar in order to retrieve a given 4-vector component for a given particle. It is only the memory access classes that, internally, are able to decode the memory layouts of a raw buffer: this is done by interpreting the buffers as one-dimensional C-style array or casting them as multi-dimensional C-style arrays. For most data buffers, a Structure-Of-Array (SOA) or an Array-Of-Structure-Of-Array (AOSOA) layout is chosen, where the data items of a given type for different events are contiguous: as explained in detail in the vCHEP2021 proceedings bib:vchep2021 , this is absolutely essential for SIMD processing on vector CPUs, and it is also beneficial – but not at all a strict requirement – on GPUs to improve performance through coalesced data access. In some cases, such as couplings, the floating point raw arrays include the real and imaginary components of complex data, and it is the responsibility of the data access classes to return outputs with the API of a complex number data type. The case of CPU code includes the additional complication of returning outputs whose data types fptype_sv and cxtype_sv can be scalar (for no-SIMD C++) or SIMD vectors fptype_v and cxtype_v (through compiler vector extensions bib:vchep2021 , hereafter CVEs) of fptype or cxtype data. Concerning the use of memory access classes in other CUDACPP software components, most notably in the computational methods for helicity amplitudes like VXXXXX, VVV1P0_1 or FFV1_0, I initially chose to implement this as a template parameter: in Table 1, for instance, DeviceAccessMomenta is a template parameter M_ACCESS of VXXXXX. The idea behind the choice of using templates was to allow the flexibility of easily switching between different memory layouts at build time, to compare their performances; in retrospective, the same flexibility may have been achieved with the same memory access classes but without templates [^1] As noted in Sec. 4, eventually I did decide to remove those templates, even if it was not clear to me a priori if this would make the code easier to read and maintain, or faster to build: my main motivation was an attempt to allow some CUDA builds in some configurations of particularly complex $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes, which were otherwise crashing or taking too long. Unfortunately, my tests about the effectiveness of template removal for simplifying CUDA builds were at best inconclusive, as those builds continued to fail; however, removing templates did not have any negative impact on performance elsewhere, and I suggest that this may represent a useful cleanup to merge upstream and integrate in a future production release. .
@@
    Arithmetic operations and other computational functions. Finally, the decoupling of data access from actual calculations has probably been one of the most important aspects of the whole CUDACPP software design. Just like the memory allocation classes, also the helicity amplitude functions like VXXXXX, VVV1P0_1 or FFV1_0 have almost [^2] One notable exception is that some assumptions are still made about the memory layout of wavefunctions, which for spin 1/2 and spin 1 particles are complex 6-dimensional arrays as in the original HELAS bib:helas and ALOHA bib:aloha implementations, from which CUDACPP is derived. This detail is relevant to the kernel splitting for different Feynman diagrams, as discussed in Sec. 4. no knowledge of the memory layouts used in the multi-event data buffers. These functions also ignore whether the calculations are performed in single or double precision, as I encapsulated this choice in the fptype type definition. Even more, these functions ignore whether the arithmetic operations within them are applied to scalar values on GPUs or CPUs or to SIMD vectors on CPUs: in the latter case, this is possible because simple operators like “+” are automatically understood by the compiler as vector operations when applied to CVE vectors of fptype, and are also implemented using SIMD CVE operations in their definition for vectors of custom types, notably the vectors of cxtype complex types. This design is extremely powerful because it has made it possible to use formally the same exact lines of software in the code-generated helicity amplitude functions like VVV1P0_1 or FFV1_0. I note in passing, however, that the functions computing wavefunctions for initial and final state particles, like IXXXXX, OXXXXX and VXXXXX, or their variations for massless or beam-collinear particles, are not code-generated from a model Lagrangian, but are instead hardcoded and often required particular care and successive iterations. Two complications that I had to address for SIMD code, in particular, are the following [^3] In this context, I find it important to stress that, in my experience, achieving a robust and performant CUDACPP implementation for SIMD CPUs was far from a trivial task, and was, in fact, much more complex than achieving it for GPUs. The difficulties for implementing IXXXXX-like functions are just two examples. I also already mentioned the fact that there are very strict constraints in the memory layouts for CPU SIMD, unlike those for GPUs; as described in Ref. bib:vchep2021 , in particular, this also implied the development of wrappers and adapters for vectors cxtype_v of complex numbers, as their real and imaginary parts must be stored as an SOA using two contiguous fptype_v arrays (RRRRIIII), rather than as an AOS (RIRIRIRI). Another example is the implementation of mixed floating point mode, where I added a complex mechanism to merge/split two SIMD vectors of double into/from one SIMD vector of float at the boundary between dual amplitude and color sum calculations, in order to use the widest possible SIMD vectors in both cases and maximize efficiency. More generally, in my opinion, it is fair to say that achieving SIMD speedups through vectorization is always much more complicated than achieving speedups on GPUs. Code acceleration on GPUs, in fact, can be achieved even without lockstep processing (at the cost of some inefficiency from warp divergence), or without optimized memory layouts (at the cost of some inefficiency from the lack of coalesced data access), simply because a GPU has thousands of threads. One good feature of CUDACPP is that, by focusing on achieving C++ vectorization on SIMD CPUs, this has also improved the efficiency of the code on GPUs. : first, unlike VVV1P0_1 or FFV1_0, the IXXXXX-like functions include some if/else branching, which I reimplemented using vector masks in the SIMD case in order to fully exploit data parallel speedups (even if, from a performance point of view, this is only important for simple processes with few Feynman diagrams); second, as I performed all my functional testing with Floating Point Exception traps (FPEs) enabled in order to develop more robust code, I came across some crashes caused by the interplay of compiler optimizations, SIMD CVEs and FPE traps, which I addressed using volatile keywords and which pushed me to develop a large set of functional tests specifically for these functions.
@@
On top of the separation of the three software concerns above (data allocation, data access and arithmetic calculations), another important part of the design of the CUDACPP software was the implementation of kernel launching on GPUs and event loops on CPUs. My aim here was to keep as much as possible of the software logic and of the actual code identical for GPUs and CPUs, in order to simplify the iterative addition of new features and of bug fixes for both cases. In short, I addressed this by keeping also for SIMD CPUs the idea of a “grid” of events that are processed in one given iteration, and by explicitly subdividing this grid into SIMD event vectors and adding an explicit loop over them. Much more practically, and more importantly for the kernel splitting work described in the next sections, the link between memory allocations and the execution of the ME computational engines is provided by two different incarnations of a MatrixElementKernel (MEK) class, one for GPUs (Device) and one for CPUs (Host). A singleton instance of the appropriate MEK class is constructed and used inside check.exe for standalone tests, or inside the Bridge component that is linked with Fortran in madevent for the full MG5aMC workflow. As shown in Table 1, it is the MatrixElementKernelDevice class that internally launches the monolithic sigmaKin kernel in the current CUDACPP. This is the starting point for the enhancements presented in the following sections.
2.3 A brief note on SIMD build modes
@@
As mentioned previously, the new developments that I present in this paper are mainly optimizations that target GPUs specifically, but they are also relevant and useful for the CUDACPP implementation on vector CPUs. In this paper, in particular, I also present results from the five different SIMD build modes of CUDACPP, namely none, sse4, avx2, 512y and 512z. While these have already been described to some extent in earlier publications from our team bib:vchep2021 ; bib:zw2025 , I find it useful to take this paper as an opportunity to give a few additional and previously undocumented details.
@@
The SIMD modes of CUDACPP were initially introduced targeting Intel CPUs. My design strategy was to define separate build modes, which on every specific system would allow not only the fastest calculation compatible with the hardware specs, but also an evaluation of the speedup achieved through the SIMD implementation of CUDACPP in vectorised C++. Largely speaking: none does not use any SIMD extensions; sse4 uses SSE4.2 instructions and 128-bit xmm registers; avx2 uses AVX2 instructions and 256-bit ymm registers; 512y uses AVX512 instructions and 256-bit ymm registers; 512z uses AVX512 instructions and 512-bit zmm registers. The 512y mode may appear surprising: the rationale behind it is, quite simply, that it is around 10% faster than avx2 on CPUs with the AVX512 instruction set, but it is also faster than 512z on AVX512 CPUs with a single FMA unit (such as Intel Xeon Silver 4216 CPUs); conversely, it is slower than 512z on AVX512 CPUs with two FMA unit (such as Intel Xeon Gold 6326 CPUs). This was first reported for the CUDACPP software in Ref. bib:ichep2022 and has been confirmed in many tests since then.
@@
    The CPU instruction set is configured via the -march flag of the compiler, in the cudacpp.mk makefile. First and foremost, this must be well-defined because an application compiled for a given instruction set would crash if it is executed on a hardware that does not support it (for instance, if it is compiled for AVX512 but executed on a host that does not support it). While many different levels and sublevels of vectorization have been defined over time, such as AVX, AVX2, SSE4.1 or SSE4.2, the choice I made was to only support a very limited subset of these in CUDACPP. In particular: none uses -march=x86-64, which disables all SIMD extensions; sse4 uses -march=nehalem, which supports SSE4.2 extensions with 128-bit xmm registers; avx2 uses -march=haswell, which supports AVX2 extensions with 256-bit ymm registers; 512y and 512z both use -march=skylake-avx512, which supports a specific set of AVX512 extensions with 512-bit zmm registers. In this context, I chose not to foresee a build mode based on -march=native, which would offer the greatest set of hardware features available on a CPU, but is somewhat poorly defined. This is similar to what is done in some LHC experiments: LHCb, for instance, uses well-defined software architecture builds based on x86_64 Micro-Architecture Feature Levels bib:phoronix such as x86_64-v2. In retrospective, indeed a better choice would have been to configure -march based on x86_64 Feature Levels, but these were not yet widely used when I developed the SIMD implementation of CUDACPP in 2020.
@@
    The SIMD vector length is the core of the CUDACPP SIMD implementation, where it is encapsulated by a preprocessor variable MGONGPU_CPPSIMD and an equivalent C++ variable neppV. As described more in detail in Ref. bib:vchep2021 , in fact, the basic strategy of the CUDACPP SIMD implementation is the definition, using compiler vector extensions, of vectors fptype_v of neppV floating point variables of type fptype. The five build modes are defined as follows: none uses neppV=1, sse4 uses neppV=2 and neppV=4 in double and single precision, respectively; avx2 and 512y both use neppV=4 and neppV=8 in double and single precision; 512z uses neppV=8 and neppV=16 in double and single precision.

### v3

@@ 2.2 A closer look at the internals of the ME engine
In the production version of the CUDACPP plugin before any kernel splitting changes (hereafter “ihel0”), v1.00.02 in MG5aMC v3.6.3, the ME computational engine on GPUs is essentially a single and very large monolithic GPU kernel, named sigmaKin. This performs all of the relevant operations that are needed to compute the output ME from the input momenta for each event (one notable exception being the calculation of other couplings that depend on $\alpha_{s}$, which is delegated to a separate kernel). Very schematically, this is represented in the pseudo-code in Table 1. In particular, sigmaKin performs the following operations for each event: it keeps a running sum of the ME over helicities; it loops on all possible helicities of the external particles (or, more precisely, only on a pre-determined set of “good” helicities whose contribution is non-zero); it adds to the running sum of the ME the contribution from each helicity, by calling a device function calculate_wavefunctions which internally involves for each QCD color flow the calculation of particle and propagator wavefunctions and that of dual amplitudes, followed by the sum over all color flows using a quadratic form based on a color matrix; after the end of the helicity loop, sigmaKin randomly draws a helicity and a color for the generated event based on the ME contributions from individual helicities and colors. The calculation of helicity amplitudes for Feynman diagrams, in particular, is described in detail in Ref. bib:vchep2021 (see Fig. 1 therein); complementary details are also given in Ref. bib:zw2025 . As shown schematically in Table 1, this involves three types of elements to compute: the wavefunctions of external (initial/final state) particles, via functions like VXXXXX; the wavefunctions of internal propagators, e.g. via VVV1P0_1; the dual amplitudes for a given color flow (“jamps”), e.g. via FFV1_0. The names and roles of these functions in CUDACPP are the same as in the original HELAS bib:helas and ALOHA bib:aloha implementations.
@@
    Memory layout and data access. The specific layout chosen for storing data inside the raw memory buffers allocated in the previous step is encapsulated in a separate set of memory access classes. All these classes have methods named kernelAccess, or some variation of this. On the GPU, these methods take as input the fptype* pointer associated to a raw memory buffer for nevt events, and they return as output a data item for the single event ievt processed by the GPU thread where the code is executed, typically indexed by the identifier of the GPU thread, blockDim.x*blockIdx.x+threadIdx.x. In some cases, these methods have additional parameters: method kernelAccessIp4IparConst in the DeviceAccessMomenta class, for instance, takes two additional parameters ip4 and ipar in order to retrieve a given 4-vector component for a given particle. It is only the memory access classes that, internally, are able to decode the memory layouts of a raw buffer: this is done by interpreting the buffers as one-dimensional C-style array or casting them as multi-dimensional C-style arrays. For most data buffers, a Structure-Of-Array (SOA) or an Array-Of-Structure-Of-Array (AOSOA) layout is chosen, where the data items of a given type for different events are contiguous: as explained in detail in the vCHEP2021 proceedings bib:vchep2021 , this is absolutely essential for SIMD processing on vector CPUs, and it is also beneficial – but not at all a strict requirement – on GPUs to improve performance through coalesced memory access. In some cases, such as couplings, the floating point raw arrays include the real and imaginary components of complex data, and it is the responsibility of the data access classes to return outputs with the API of a complex number data type. The case of CPU code includes the additional complication of returning outputs whose data types fptype_sv and cxtype_sv can be scalar (for no-SIMD C++) or SIMD vectors fptype_v and cxtype_v (through compiler vector extensions bib:vchep2021 , hereafter CVEs) of fptype or cxtype data. Concerning the use of memory access classes in other CUDACPP software components, notably the methods for computing helicity amplitudes like VXXXXX, VVV1P0_1 or FFV1_0, I initially chose to implement this via templates: DeviceAccessMomenta in Table 1, for instance, is a template parameter M_ACCESS of VXXXXX. The idea behind the choice of using templates was to allow the flexibility of easily switching between different memory layouts at build time, to compare their performances; in retrospective, the same flexibility may have been achieved with the same memory access classes but without templates [^1] As noted in Sec. 4.4, eventually I did decide to remove those templates, even if it was not clear to me a priori if this would make the code easier to read and maintain, or faster to build: my main motivation was an attempt to allow some CUDA builds in some configurations of particularly complex $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes, which were otherwise crashing or taking too long. Unfortunately, my tests about the effectiveness of template removal for simplifying CUDA builds were at best inconclusive, as those builds continued to fail; however, removing templates did not have any negative impact on performance elsewhere, and I suggest that this may represent a useful cleanup to merge upstream and integrate in a future production release. .
@@
    Arithmetic operations and other computational functions. Finally, the decoupling of data access from actual calculations has probably been one of the most important aspects of the whole CUDACPP software design. Just like the memory allocation classes, also the helicity amplitude functions like VXXXXX, VVV1P0_1 or FFV1_0 have almost [^2] One notable exception is that some assumptions are still made about the memory layout of wavefunctions, which for spin 1/2 and spin 1 particles are complex 6-dimensional arrays as in the original HELAS bib:helas and ALOHA bib:aloha implementations, from which CUDACPP is derived. This detail is relevant to the kernel splitting for different Feynman diagrams, as noted in Sec. 4.2. no knowledge of the memory layouts used in the multi-event data buffers. These functions also ignore whether the calculations are performed in single or double precision, as I encapsulated this choice in the fptype type definition. Even more, these functions ignore whether the arithmetic operations within them are applied to scalar values on GPUs or CPUs or to SIMD vectors on CPUs: in the latter case, this is possible because simple operators like “+” are automatically understood by the compiler as vector operations when applied to CVE vectors of fptype, and are also implemented using SIMD CVE operations in their definition for vectors of custom types, notably the vectors of cxtype complex types. This design is extremely powerful because it has made it possible to use formally the same exact lines of software in the code-generated helicity amplitude functions like VVV1P0_1 or FFV1_0. I note in passing, however, that the functions computing wavefunctions for initial and final state particles, like IXXXXX, OXXXXX and VXXXXX, or their variations for massless or beam-collinear particles, are not code-generated from a model Lagrangian, but are instead hardcoded and often required particular care and successive iterations. Two complications that I had to address for SIMD code, in particular, are the following [^3] In this context, I find it important to stress that, in my experience, achieving a robust and performant CUDACPP implementation for SIMD CPUs was far from a trivial task, and was, in fact, much more complex than achieving it for GPUs. The difficulties for implementing IXXXXX-like functions are just two examples. I also already mentioned the fact that there are very strict constraints in the memory layouts for CPU SIMD, unlike those for GPUs; as described in Ref. bib:vchep2021 , in particular, this also implied the development of wrappers and adapters for vectors cxtype_v of complex numbers, as their real and imaginary parts must be stored as an SOA using two contiguous fptype_v arrays (RRRRIIII), rather than as an AOS (RIRIRIRI). Another example is the implementation of mixed floating point mode, where I added a complex mechanism to merge/split two SIMD vectors of double into/from one SIMD vector of float at the boundary between dual amplitude and color sum calculations, in order to use the widest possible SIMD vectors in both cases and maximize efficiency (see Sec. 3.5). More generally, in my opinion, it is fair to say that achieving SIMD speedups through vectorization is always much more complicated than achieving speedups on GPUs. Code acceleration on GPUs, in fact, can be achieved even without lockstep processing (at the cost of some inefficiency from warp divergence), or without optimized memory layouts (at the cost of some inefficiency from the lack of coalesced data access), simply because a GPU has thousands of threads. One good feature of CUDACPP is that, by focusing on achieving C++ vectorization on SIMD CPUs, this has also improved the efficiency of the code on GPUs. : first, unlike VVV1P0_1 or FFV1_0, the IXXXXX-like functions include some if/else branching, which I reimplemented using vector masks in the SIMD case in order to fully exploit data parallel speedups (even if, from a performance point of view, this is only important for simple processes with few Feynman diagrams); second, as I performed all my functional testing with Floating Point Exception traps (FPEs) enabled in order to develop more robust code, I came across some crashes caused by the interplay of compiler optimizations, SIMD CVEs and FPE traps, which I addressed using volatile keywords and which pushed me to develop a large set of functional tests specifically for these functions.
@@
On top of the separation of the three software concerns above (data allocation, data access and arithmetic calculations), another important part of the design of the CUDACPP software was the implementation of kernel launching on GPUs and event loops on CPUs. My aim here was to keep as much as possible of the software logic and of the actual code identical for GPUs and CPUs, in order to simplify the iterative addition of new features and of bug fixes for both cases. In short, I addressed this by keeping also for SIMD CPUs the idea of a “grid” of events that are processed in one given iteration, and by explicitly subdividing this grid into SIMD event vectors and adding an explicit loop over them. Much more practically, and more importantly for the kernel splitting work described in the following, the link between memory allocations and the execution of the ME computational engines is provided by two different incarnations of a MatrixElementKernel (MEK) class, one for GPUs (Device) and one for CPUs (Host). A singleton instance of the appropriate MEK class is constructed and used inside check.exe for standalone tests, or inside the Bridge component that is linked with Fortran in madevent for the full MG5aMC workflow. As shown in Table 1, it is the MatrixElementKernelDevice class that internally launches the monolithic sigmaKin kernel in the current CUDACPP. This is the starting point for the enhancements presented in the following sections.
2.3 A few details on SIMD build modes
@@
As mentioned previously, the new developments that I present in this paper are mainly optimizations that target GPUs specifically, but they are also relevant and useful for the CUDACPP implementation on vector CPUs. In the following, in particular, I will also show results from the five different SIMD build modes of CUDACPP, namely none, sse4, avx2, 512y and 512z. While these have already been described to some extent in earlier publications from our team bib:vchep2021 ; bib:zw2025 , I find it useful to take this paper as an opportunity to give a few additional and previously undocumented details. This is also relevant to the further optimizations of color sums on vector CPUs that are described in Sec. 3.5.
@@
The SIMD modes of CUDACPP were initially introduced targeting Intel CPUs. My design strategy was to define separate build modes, which on every specific system would allow not only the fastest calculation compatible with the hardware specs, but also an evaluation of the speedup achieved through the SIMD implementation of CUDACPP in vectorized C++. Largely speaking: none does not explicitly use any SIMD extensions; sse4 uses SSE4.2 instructions and 128-bit xmm registers; avx2 uses AVX2 instructions and 256-bit ymm registers; 512y uses AVX512 instructions and 256-bit ymm registers; 512z uses AVX512 instructions and 512-bit zmm registers.
The 512y mode may appear surprising. The rationale behind it is, quite simply, that it is around 10% faster than avx2 on CPUs with the AVX512 instruction set: as can be seen by disassembling CUDACPP objects with the objdump tool (see slide 17 in Ref. bib:vchep2021video ), this is probably because the compiler chooses to process a few operations using AVX512 instructions rather than AVX2 instructions. In addition, 512y is also faster than 512z on AVX512 CPUs with a single FMA unit (such as Intel Xeon Silver 4216 CPUs), making it the preferred CUDACPP build mode on this large family of CPUs. Conversely, 512y is slower than 512z, and the latter should be preferred, on AVX512 CPUs with two FMA units (such as Intel Xeon Gold 6326 CPUs): this was first reported for the CUDACPP software in Ref. bib:ichep2022 and has been confirmed in many tests since then.
@@
    CPU instruction set. This is configured via the compiler flag -march in the cudacpp.mk makefile. First and foremost, this must be well-defined because an application built for a given instruction set would crash if it is executed on a hardware that does not support it (for instance, if it is compiled for AVX512 but executed on a host that does not support it). While many different levels and sublevels of vectorization have been defined over time, such as AVX, AVX2, SSE4.1 or SSE4.2, the choice I made was to only support a very limited subset of these in CUDACPP. In particular: none uses -march=x86-64, which is meant to disable all SIMD extensions (but does not completely achieve this goal, as discussed in Sec. 3.5); sse4 uses -march=nehalem, which supports SSE4.2 extensions with 128-bit xmm registers; avx2 uses -march=haswell, which supports AVX2 extensions with 256-bit ymm registers; 512y and 512z both use -march=skylake-avx512, which supports a specific set of AVX512 extensions with 512-bit zmm registers. In this context, I chose not to foresee a build mode based on -march=native, which would offer the greatest set of hardware features available on a CPU, but is somewhat poorly defined. This is similar to what is done in some LHC experiments: LHCb, for instance, uses well-defined software architecture builds based on x86_64 Micro-Architecture Feature Levels bib:phoronix such as x86_64-v2. In retrospective, indeed a better choice would have been to configure -march based on x86_64 Feature Levels, but these were not yet widely used when I developed the SIMD implementation of CUDACPP in 2020.
@@
    SIMD vector length. This is the core of the CUDACPP SIMD implementation, where it is encapsulated by a preprocessor variable MGONGPU_CPPSIMD and an equivalent C++ variable neppV. As described more in detail in Ref. bib:vchep2021 , in fact, the basic strategy of the CUDACPP SIMD implementation is the definition, using compiler vector extensions, of vectors fptype_v of neppV floating point variables of type fptype. The five build modes are defined as follows: none uses neppV=1, sse4 uses neppV=2 and neppV=4 in double and single precision, respectively; avx2 and 512y both use neppV=4 and neppV=8 in double and single precision (i.e. this is the same source code compiled with different flags); 512z uses neppV=8 and neppV=16 in double and single precision.